OpenStudy (anonymous):

I'll medal to who ever assists me! <3 What is the 43rd term of an arithmetic sequence with a rate of increase of -6 and a11 = 12?} A.) -174 B.) -176 C.) -180 D.) -186 E.) -240 Arigato! ^-^

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[a_{11}=12\] right?

OpenStudy (misty1212):

so \[a_{12}=12-6\\ a_{13}=12-2\times 6\\ a_{14}=12-3\times 6\\ a_{15}=12-4\times 6\] etc

OpenStudy (misty1212):

reason i am writing it out is because it it confusing with the subscript do you get what the \(43\) term would be?

OpenStudy (misty1212):

did i lose you?

OpenStudy (anonymous):

yes you did hahaha @misty1212